# AgentA/B: Automated and Scalable Web A/B Testing with Interactive LLM Agents

### v2

@@ Abstract.
A/B testing experiment is a widely adopted method for evaluating UI/UX design decisions in modern web applications. Yet, traditional A/B testing remains constrained by its dependence on the large-scale and live traffic of human participants, and the long time of waiting for the testing result. Through formative interviews with six experienced industry practitioners, we identified critical bottlenecks in current A/B testing workflows. In response, we present AgentA/B, a novel system that leverages Large Language Model-based autonomous agents (LLM Agents) to automatically simulate user interaction behaviors with real webpages. AgentA/B enables scalable deployment of LLM agents with diverse personas, each capable of navigating the dynamic webpage and interactively executing multi-step interactions like search, clicking, filtering, and purchasing. In a demonstrative controlled experiment, we employ AgentA/B to simulate a between-subject A/B testing with $1,000$ LLM agents Amazon.com, and compare agent behaviors with real human shopping behaviors at a scale. Our findings suggest AgentA/B can emulate human-like behavior patterns.
@@
Figure 1. Architecture of AgentA/B system. The system takes user input of agent and testing specifications to automatically 1) generate a large-scale of LLM agents, 2) perform agent-web interaction with the assigned web environment, and 3) conduct post-testing analysis, and return the results to the end users.
[…]
A flowchart diagram illustrates the A/B testing workflow enabled by UXAgent. On the left, a system user (e.g., UX researcher or product manager) provides two types of input: LLM Agent specifications (e.g., number of agents, demographic distributions, attributes) and A/B Testing configuration (e.g., feature designs, agent intentions, metrics). These inputs lead to two parallel processes: LLM Agent Generation (Step 1) and Testing Condition Preparation (Step 2). The generated agents are then assigned to either a control or treatment group (Step 3), where they interact with their designated web environments. The control group uses the baseline interface, while the treatment group uses a modified (new) design. Simulation results are collected (Step 4) and passed to the Post-Testing Analysis module (Step 5), which returns A/B Testing Results (e.g., session outcomes, average actions, failure rates) back to the user. The visual includes labeled arrows, icons representing agents and interfaces, and text boxes describing each component.
@@ 1. Introduction
In the fast-paced online economy, user-facing web design decisions are central to user engagement and satisfaction. A/B testing–also known as online controlled experimentation–has become a cornerstone methodology for rapidly evaluating user interface designs and new system features (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015). A/B testing involves comparing two or more variants of a webpage (could be slight design changes or entirely new functionality) with real users randomly routed to the controlled group and the treatment (or experiment) group to evaluate the impact of design decisions on user engagements (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021). Major technology companies, including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), build their own A/B testing infrastructure and conduct hundreds of A/B tests to optimize their products and marketing strategies.
Despite being widely adopted, traditional A/B testing is inherently constrained by a number of participant-related challenges. First, online A/b testing requires a large sample size of users to achieve statistically significant results (often hundreds of thousands), but live user traffic is often expensive and resource-intensive, which may be especially problematic for niche markets or newer websites with low traffic (Kohavi et al., 2013; Xu et al., 2015). Second, user feedback cycles are inherently slow because a typical A/B test involves complex procedures such as experiment design, feature implementation, traffic allocation, result monitoring, and post-testing analysis (Fabijan et al., 2018), among which the result monitoring step alone may take weeks to complete. Third, the complexity of traffic coordination, engineering effort, and other administrative configurations means that the design teams can rarely test more than a handful of ideas throughout the design life-cycle. In practice, this means that many promising design variants never get tested due to operational constraints.
Recent advances in Large Language Model powered agents (LLM agents) offer a promising avenue to mitigate these limitations (Chen et al., 2025). LLM agents have demonstrated remarkable capabilities to mimic human decision-making, generate plausible behavior sequences, and reason across a range of domains, including healthcare (Li et al., 2024), software engineering (Luo et al., 2023), social science (Park et al., 2023; Hua et al., 2023), and autonomous agents (Wu et al., 2023; Yao et al., 2023). In particular, researchers have started to deploy LLM agents in web environments for tasks such as shopping (Yao et al., 2022), service booking (Zhou et al., 2023), and multimodal search (Koh et al., 2024). However, the vast majority of these systems operate in a single-agent session or work only on sandbox web environments, which often fail to capture the dynamic and diverse nature of the large-scale user interaction sessions on real-world websites.
In the context of A/B testing, we hypothesize LLM agents can help evaluate multiple design variants of web designs by role-playing real users’ interactions with an interactive web environment. This possibility could enable website designers and product managers to rapidly and risk-freely piloting many design candidates without worrying about limited user traffic. Our position is that LLM agents should not replace real user testings in a feature design project lifecycle; instead, it could complement the existing but limited piloting sessions for design optimization.
Our project began with a formative study with six experienced professionals who conduct A/B testing in their day-to-day workflow and extracted a current A/B testing workflow (Fig 2) and their challenges: 1) the lack of lightweight prototyping tools for early hypothesis testing, 2) the scarcity and contestation of user traffic as an experimental resource, and 3) the long turnaround times for user feedback that discourage design exploration. Based on this, we then developed AgentA/B, an LLM agent-based end-to-end A/B testing system, with the following four functions (Fig. 1): 1) LLM agent persona generation, 2) A/B testing condition preparation, 3) iterative and automated agent-web interaction, and 3) post-testing data analysis. Our system can plug and play with any existing LLM agent systems (e.g., Claude Computer-use agent or ReAct (Yao et al., 2023)) to generate hundreds of thousands of virtual users, each with a generated user persona, and these virtual users can interactively observe and operate on real webpages. For each agent-web interaction session (Fig. 3), our system has a web environment parsing module to preprocess dynamic web pages into structured, semantically meaningful JSON observations for LLM agents to consume, and has an action execution module to translate the LLM agent generated next action (e.g., “solar filter for telescope”) into a web operation (“click on solar filter” and “click first product”). The system monitors the LLM-agent reasoning logic and the agent behavior trajectories as a virtual user interaction session and automatically generates post-testing analysis results for users to analyze (e.g., t-test of purchased price in the two groups).
We use Amazon.com as a case study for automated A/B testing, where we generated 100,000 virtual customer personas and randomly selected $1,000$ to generate an LLM agent-based virtual customer using each of the personas. The design to be evaluated is whether the webpage shows all filter options in the left panel (control) or only shows a reduced list of options (treatment). The results show that LLM agents in the treatment group conducted more purchase and filter actions compared to those in the control group, which aligns with the direction of effects observed in the parallel human study. We also compared the behavior of LLM agents with aggregated results from 1 million human users and found that LLM agents were more goal-directed in their interactions, resulting in a shorter action length.
In summary, this paper implements AgentA/B, a system that deploys LLM-based agents into A/B testing conditions to interact with live websites to simulate realistic user interactions. Our case study of a 1,000 agent between-subject A/B testing on Amazon.com shows promising results of using the AgentA/B to conduct an automated and scalable Web A/B testing.
@@ 2.1. Limitations of Traditional A/B Testing in Practice
Figure 2. The workflow of web A/B testing and three challenges reported from formative study: 1) the cost and difficulty of securing considerable user traffic for significant results, 2) the whole A/B testing period can span across weeks and months, and 3) limited testing opportunites.
[…]
A timeline diagram illustrates the standard A/B testing workflow and highlights three key challenges. The timeline includes seven sequential phases: Feature Ideation (new UI layout or functionalities), A/B Testing Experiment Design (testing configuration and implementation), Data Monitoring (collect user interactions and rationales), Feature Deployment (final design decisions), Post-Testing Analysis (behavior patterns, failure cases, outcomes), Traffic Allocation (real user traffic toward variants), and Team Alignment (decision maker approval or rejection). The A/B Testing Period is marked across the Data Monitoring and Traffic Allocation stages. Three challenges are highlighted in red: (1) Scalable recruitment difficulties, (2) Limited testing opportunities, and (3) Slow user feedback loop. Each stage is connected along a central purple timeline.
@@ 2.3. User Behavior Simulation: From Cognitive Models to LLM Agents
With the rise of large language models (LLMs), researchers have begun to explore how LLMs can simulate human-like behaviors across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024). Park et al. (2023) built a generative society of agents simulating daily social behavior in a simulated town. Horton (2023) evaluated whether LLMs could replicate outcomes from real behavioral experiments on prosociality. Furthermore, Lu et al. (2025a) demonstrates that fine-tuning large language models (LLMs) on real-world behavioral data significantly enhances their ability to generate accurate user actions. These systems suggest that LLMs can produce high-fidelity behavioral simulations across diverse tasks. Our work extends this direction into the visually rich, dynamic domain of live web interfaces, focusing on shopping behavior and automated A/B testing experiments.
@@ 3.2. Formative Study Findings
4. AgentA/B: Automated A/B Testing on the Web with LLM Agents
@@ 4.1. System Overview and Pipeline
The post-test analysis module outputs summary statistics such as actions per session, session duration (in steps and time), and purchase completion rate. Researchers can also examine detailed behaviors (e.g., search or click filter usage) and compare them across A/B condition variants. The system supports stratified analysis by agent demographics or personas to identify subgroup differences. It automatically computes effect sizes (e.g., absolute/relative changes in conversion-like behaviors) to aid interpretation. For instance, when testing redesigned filters, the system can reveal whether agents refined searches more, completed tasks faster, and purchased more, which offers early insights on usability and adoption risks before live deployment.
Finally, AgentA/B maintains compatibility with external analytics pipelines. The result logs (in .JSON and .XSL) are exported in a format compatible with common data science tools, enabling downstream statistical modeling, significance testing, or integration with traditional A/B testing dashboards. Users can use AgentA/B to analyze thousands of simulated interaction session results into actionable insights for interface evaluation.
@@
Figure 3. One action prediction iteration of the automated web testing in AgentA/B with an LLM agent. (1) An Agent Profiling Module maintains a comprehensive agent description with an LLM-generated persona, user-specified intention, and the action history of the current session. In the meantime, (2) the Environment Parsing Module parses the webpage into structured web representation and action spaces. (3) All the information is fed into the LLM Agent for the next action prediction, (4) which will be executed by the Action Execution Module in the web environment to drive the next iteration.
[…]
A system diagram shows how a language model-based agent interacts with a web environment. The process begins with a web interface (Control/Treatment) feeding into an Environment Parsing Module, which extracts the webpage as a JSON structure and defines an action space (e.g., search, filter, purchase, product_click, terminate). This information, along with a predefined persona and intention, is sent to the LLM Agent (Action Prediction), which outputs the next action (e.g., search “solar filter for telescope”). This action is executed through the Action Execution Module, which sends it back to the Web Environment, completing the interaction loop.
@@ 4.2. Agent-Environment Interaction Architecture
Figure 4. The two design variants of the left filter panel on Amazon.com for the A/B testing case study with AgentA/B. In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the users’ search query (e.g., “solar filter for telescope”).
[…]
A side-by-side comparison of two Amazon search results pages for the query ”solar filter for telescope.” Both pages show similar products but have differences in their filter options on the left sidebar. The left image displays additional filter categories under ”Department” such as ”Telescope Refractors,” ”Safety Goggles & Glasses,” and ”Telescope Eyepieces,” which are not present in the right image. Both images include nested subcategories under ”Camera & Photo Filters” and identical product listings on the right side of the page. This comparison highlights subtle differences in the user interface that could affect user navigation or filtering behavior.
@@
The next action predicted by the LLM agent is translated into browser commands by the Action Execution Module. Actions are expressed in a structured format that can reference DOM elements or logical operations, such as Click_product(3), Click_filter_option(Brand: Sony), Search(”Wireless earbuds”), or Purchase. The execution module parses the action and performs the corresponding interaction on the live webpage. In some cases, the web execution is not guaranteed to succeed, as real-world pages are prone to dynamic content loading and modal interruptions. Therefore, the execution module incorporates built-in fault detection and recovery logic. If an action fails due to a missing selector or DOM mismatch, the system attempts fallback options such as retrying, scrolling into view, or re-parsing the page. Each successful or failed execution updates the environment and starts the next iteration of the loop.
@@ 5.1. Method: A/B Testing Scenario
We designed a simulated A/B testing scenario focused on the left-side filtering panel of Amazon’s shopping interface, a common area for UI experimentation in e-commerce. The goal was to evaluate how variations in filter configurations influence user behavior.
{booktabs}
colspec=lccc Control Condition
@@
Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and randomly sampled 1,000 to simulate individual shopping sessions. Each agent was initialized with a persona profile and a shopping goal (e.g., “find a budget smart speaker under \$40 with strong customer reviews”). The allowed action space includes typical e-commerce flows: search, filter use, product examination, and cart actions. Persona generation followed the methodology in Chen et al. (2024) with LLM agent’s demographics (age, income, occupation), preferences, and shopping goals.
In the case study, our AgentA/B was implemented with the Claude 3.5 Sonnet model as the LLM backend to support agent generation, testing preparation, automated agent-web interaction, and post-testing analysis. The AgentA/B environment was executed on a distributed cluster of 16 high-memory compute nodes, with each node controlling a Selenium-driven Chrome instance running in headless mode. Each session was capped at 20 actions. Sessions ended either with task success (completion or no further predicted actions) or failure (e.g., looping behavior). We logged full action traces, metadata (e.g., duration, outcome), and agent rationales where applicable.
@@ 5.2. Finding: Alignment with Human Behavior
To benchmark our simulation system in aligning with human behavior, we conducted a simulated A/B testing case study using AgentA/B and also have access to the results of online A/B testing results with real human subjects using identical task conditions. Table 5.1 presents the aggregated actions and statistics from sessions with LLM agents and with human participants.
Human participants engaged in longer, more exploratory interactions than agents, averaging nearly twice as many actions per session (15.96 vs. 6.05). They also clicked more products (6.96 vs. 1.87) and performed more searches (6.40 vs. 1.42). Despite these differences in interaction style, humans and agents showed similar purchase rates (0.62 vs. 0.81) and comparable use of filter options (0.33 vs. 0.58).

### v3

@@ Abstract.
A/B testing is central to UI/UX design, yet our formative study with six industry practitioners revealed that it is slowed by scarce user traffic, long run times, and high operational cost. To address these challenges, we introduce Agent A/B, an end-to-end system that deploys large language model (LLM) agents with structured personas to interact with live webpages and generate scalable and behavioral evidence before launch. In a case study on Amazon.com, Agent A/B simulated a between-subjects A/B test of filter panel designs with 1,000 agents and found that the reduced filter list produced more purchases (414 vs. 403), reproducing directional outcomes also observed in a parallel large-scale human experiment. Results further suggest that agent-based simulations can detect interface-sensitive behavioral differences and surface subgroup patterns while offering faster, lower-risk insights. We position Agent A/B as a complement to human testing, enabling earlier prototyping, pre-deployment validation, and hypothesis-driven UX evaluation.
@@
Figure 1. Architecture of Agent A/B system. The system takes user input of agent and testing specifications to automatically 1) generate a large-scale of LLM agents, 2) prepare the testing conditions (e.g., traffic allocation), 3) perform agent-web interaction with the assigned web environment, 4) monitor and record agent behaviors as well as web interactions, and 5) conduct post-testing analysis so that the results will be returned to the end users.:
@@ 1. Introduction
A/B testing—also known as online controlled experimentation—has become a cornerstone methodology for evaluating design decisions in modern web applications (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015). By comparing alternative interface designs or feature functionalities with real users randomly assigned to control and treatment groups, A/B testing provides rigorous, data-driven evidence for user experience optimization (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021). Major technology companies,including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), rely on hundreds of such tests each year to refine features, improve engagement, and guide strategic design choices. Despite its ubiquity, however, traditional A/B testing is often slowed by practical constraints, including heavy reliance on live user traffic, long turnaround times for statistically reliable results, and significant engineering and organizational overhead (Kohavi et al., 2013; Xu et al., 2015; Fabijan et al., 2018).
To better understand these challenges, we conducted a formative study with six industry practitioners who regularly run A/B tests in their day-to-day workflows. Their accounts highlight three recurring bottlenecks (Fig 2). First, there is a lack of lightweight methods for piloting design hypotheses before committing to large-scale traffic allocation. Second, user traffic itself is a scarce and contested resource, particularly for niche products or teams with competing priorities. Third, feedback cycles are slow: weeks may pass between experiment setup and actionable insights, discouraging exploration of alternative designs. Collectively, these constraints mean that many promising ideas are never tested, and early-stage prototyping rarely benefits from rigorous, behavioral-data grounded evaluation.
Recent advances in large language model (LLM)–based agents offer a promising way to address these limitations (Chen et al., 2025). LLM agents have demonstrated the ability to role-play user personas, generate plausible sequences of decisions, and perform multi-step interactions in domains such as e-commerce (Yao et al., 2022), booking (Zhou et al., 2023), and information search (Koh et al., 2024). Yet most prior work has focused on single-session tasks or constrained sandbox environments (Li et al., 2024; Luo et al., 2023; Park et al., 2023; Hua et al., 2023; Wu et al., 2023; Yao et al., 2023), limiting their relevance for evaluating the dynamic and diverse behaviors seen on real-world websites. What remains missing is a scalable framework that can deploy large numbers of persona-driven agents into live web environments, allowing designers to test interface variants and obtain early behavioral signals before committing real user traffic.
We address this gap by proposing Agent A/B (Fig. 1), an end-to-end system for simulating large-scale online experiments using LLM-based agents. The system is designed to be interoperable, able to plug and play with existing agent stacks (e.g., Claude computer-use agents or ReAct (Yao et al., 2023)). As shown in Fig. 1, Agent A/B integrates five coordinated components, spanning agent specification, testing configuration, live agent–web interaction, behavioral monitoring, and automated post-testing analysis. At the core of this pipeline lies the LLM agent sub-system (Fig. 3), which iteratively parses dynamic webpages into structured representations, reasons about the next action given its persona and intention, and grounds that action into concrete DOM operations on live sites. This design enables thousands of agents to run distributed sessions at scale, producing rich traces and statistical comparisons without manual intervention.
Our position is that LLM agents should not replace real user testing; instead, they complement limited piloting sessions by offering earlier, faster, and lower-risk insights for design optimization.
We demonstrate the system through a case study on Amazon.com, where we simulated a between-subjects A/B test comparing the existing full filter panel design against a reduced filter panel novel design powered by a similarity-based ranking algorithm. Using 1,000 agents (500 per condition), we observed that the reduced filter list led to more purchases (414 vs. 403), a modest but statistically reliable increase. Average spending showed a small upward trend, though not statistically significant, and exploratory subgroup analyses suggested stronger effects for male and older personas. Importantly, these agent-based outcomes aligned directionally with the results of a parallel large-scale human A/B experiment, lending confidence to the system’s ability to surface meaningful behavioral signals.
In summary, this work contributes:
  1. (1)
    Agent A/B, a system that deploys LLM-based agents to conduct scalable, persona-driven A/B testing on live webpages.
  2. (2)
    Empirical evidence from a case study on Amazon.com showing alignment between agent simulations and large-scale human A/B results.
  3. (3)
    Future design direction, demonstrating how agent-based simulations can extend A/B testing toward earlier prototyping, inclusive piloting, and automated design exploration.
@@ 2.1. Limitations of Traditional A/B Testing in Practice
Figure 2. The workflow of web A/B testing and three challenges reported from formative study: 1) the cost and difficulty of securing considerable user traffic for significant results, 2) the whole A/B testing period can span across weeks and months, and 3) limited testing opportunites. :
@@ 2.3. User Behavior Simulation: From Cognitive Models to LLM Agents
With the rise of large language models (LLMs), researchers have begun to explore how LLMs can simulate human behaviors across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024). Park et al. (2023) built a generative society of agents simulating daily social behavior in a simulated town. Horton (2023) evaluated whether LLMs could replicate outcomes from real behavioral experiments on prosociality. Furthermore, Lu et al. (2025a) demonstrates that fine-tuning large language models (LLMs) on real-world behavioral data significantly enhances their ability to generate accurate user actions. These systems suggest that LLMs can produce high-fidelity behavioral simulations across diverse tasks. Our work extends this direction into the visually rich, dynamic domain of live web interfaces, focusing on shopping behavior and automated A/B testing experiments.
@@ 3.2. Formative Study Findings
4. Agent A/B
@@ 4.1. System Overview and Pipeline
The post-test analysis module outputs summary statistics such as actions per session, session duration (in steps and time), and purchase completion rate. Researchers can also examine detailed behaviors (e.g., search or click filter usage) and compare them across A/B condition variants. The system supports stratified analysis by agent demographics or personas to identify subgroup differences. For instance, when testing redesigned filters, the system can reveal whether agents refined searches more, completed tasks faster, and purchased more, which offers early insights on usability and adoption risks before live deployment.
Finally, Agent A/B maintains compatibility with external analytics pipelines. The result logs (in .JSON and .XLS) are exported in a format compatible with common data science tools, enabling downstream statistical modeling, significance testing, or integration with traditional A/B testing dashboards. Users can use Agent A/B to analyze thousands of simulated interaction session results into actionable insights for interface evaluation.
@@
Figure 3. One action prediction iteration of the automated web testing in Agent A/B with an LLM agent. (1) An Agent Profiling Module maintains a comprehensive agent description with an LLM-generated persona, user-specified intention, and the action history of the current session. In the meantime, (2) the Environment Parsing Module parses the webpage into structured web representation and action spaces. (3) All the information is fed into the LLM Agent for the next action prediction, (4) which will be executed by the Action Execution Module in the web environment to drive the next iteration. :
@@ 4.2. Agent-Environment Interaction Architecture
Figure 4. The two design variants of the left filter panel on Amazon.com for the A/B testing case study with Agent A/B. In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the users’ search query (e.g., “solar filter for telescope”).:
@@
The next action predicted by the LLM agent is translated into browser commands by the Action Execution Module. Actions are expressed in a structured format that can reference DOM elements or logical operations, such as Click_product(3), Click_filter_option(Brand: Sony), Search(”Wireless earbuds”), or Purchase. The execution module parses the action and performs the corresponding interaction on the live webpage. In some cases, the web execution is not guaranteed to succeed, as real-world pages are prone to dynamic content loading and modal interruptions. Therefore, the execution module incorporates built-in fault detection and recovery logic. If an action fails due to a missing selector or DOM mismatch, the system attempts fallback options such as retrying, scrolling into view, or re-parsing the page. Each execution updates the environment and starts the next iteration of the loop.
@@ 5.1. Method: A/B Testing Scenario
We designed a simulated A/B testing scenario focused on a novel feature design of the left-side filtering panel of Amazon’s shopping interface, a common area for UI experimentation in e-commerce. The goal was to evaluate how variations in filter configurations influence user behaviors such as dollar spending and product purchases.
Control Condition
@@
Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and randomly sampled 1,000 to simulate individual shopping sessions, 500 assigned into the treatment condition and the other 500 assigned into the control condition. In parallel, we also conducted the real human online A/B testing with 1 million users randomly assigned into the control condition and another 1 million users in the treatment condition. Each agent was initialized with a persona profile and a shopping task (e.g., “find a budget smart speaker under \$40 with strong customer reviews”). The allowed action space for the LLM agent includes typical e-commerce flows: search, filter use, product examination, and cart actions. Persona generation followed the methodology in Chen et al. (2024) with LLM agent’s demographics (age, income, occupation), preferences, and shopping goals.
In the case study, our Agent A/B was implemented with the Claude 3.5 Sonnet model as the LLM backend to support agent generation, testing preparation, automated agent-web interaction, and post-testing analysis. The Agent A/B environment was executed on a distributed cluster of 16 high-memory compute nodes, with each node controlling a Selenium-driven Chrome instance running in headless mode. Each session was capped at 20 actions. Sessions ended either with task success (completion or no further predicted actions) or failure (e.g., looping behavior). We logged full action traces, metadata (e.g., duration, outcome), and the agent’s rationales where applicable.
@@ 5.2. Finding: Alignment with Human Behavior
To benchmark our simulation system in aligning with human behavior, we conducted a simulated A/B testing case study using Agent A/B and also have access to the results of online A/B testing results with real human subjects using identical task conditions. Table 2 presents the aggregated actions and statistics from sessions with the 1,000 LLM agents participating and with 2M human participants [^4] Due to company policy, the human participant’s behavior data in the treatment condition can not be shared..
As shown in Table 2, human participants in the control condition engaged in longer, more exploratory interactions than agents, averaging nearly twice as many actions per session (15.96 vs. 6.05). They also clicked on more product detail pages (6.96 vs. 1.87) and performed more searches (6.40 vs. 1.42). Despite these differences in interaction style, humans and agents showed similar purchase rates (0.62 vs. 0.81) and comparable use of filter options (0.33 vs. 0.58).
@@
Figure 5. Number of LLM agents who completed a purchase under control and treatment conditions.:
(a) Average customer spending by gender groups across conditions.
(b) Average customer spending by age groups across conditions.
Figure 6. Comparison of average customer spending across conditions, broken down by gender (a) and by age groups (b).: